OpenStudy (anonymous):

Find tan θ if sec θ = square root 26/5 and sin θ < 0.

OpenStudy (anonymous):

Do you know how to draw the angle on the unit circle?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Do you know what a unit circle is?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Is your value\[\frac{ \sqrt{26} }{5 }\]or is it \[\sqrt{\frac{ 26 }{ 5 }}\]

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

@FutureMathProfessor

OpenStudy (anonymous):

What is the definintion of sec?